OpenStudy (technodynamic):

Iron has a density of 7.9 g/cm3. What is the mass of a cube of iron with the length of one side equal to 55.0 mm? Please explain step-by-step. I'll award you with a trophy too for your assistance. Thanks.

OpenStudy (anonymous):

Trust me this question is way easier then you think it is... First of all convert 55.0 mm to cm \[55mm \times \frac{ 1cm }{ 10mm }=55cm\] Second convert 'cm' to 'cm^3' to get volume, using the following equation\[V=l ^{3}\] Where 'V' is volume, 'l' is length. The reason for this equation is because to find volume you have to multiply the following: length*width*height. Since cube has all the same lengths, using the equation above is ok. So let's do it.\[V=(5.5)^{3}=166.37cm ^{3}\] Great now we have volume now we just need it to cancel out the volumes in density as follows.\[\frac{ grams }{ volume}\times volume=grams\] So the answer would be:\[Grams=\frac{ 7.9g }{ cm ^{3} }\times166.37cm ^{3}=1314.32g\]